If a + 1 = b, then b < a.
If a + 1 = b, then b > a.
Which is true and which is false?

Answer by Edwin McCravy(20060)



If we add 1 to a number, we always get a BIGGER number.

a + 1 = b  says:

If we add 1 to "a", we get "b".

So "b" is a BIGGER number than a, so we use the one
of these > or < which has the BIG side facing toward b,
and the SMALL side facing toward b.  
So,

If a + 1 = b, then b < a     <----false
If a + 1 = b, then b > a.    <----true

Edwin